# Designing Library of Skill-Agents for Hardware-Level Reusability

### 04 Mar 2024 v1

@@ 3.4.2 TR-NC and NC-TR
The TR (M=2, D=0, C=1) has two maintenance dimensions and one constraint dimension and a typical example of an object in this state is a cube sandwiched between two parallel walls. In this case, the normal direction of the wall is in the constraint dimension. The cube can move in the maintenance directions (i.e., in the directions parallel to the walls). This maintenance motion itself does not cause a state transition of the object in an infinitesimal interval, but the transition occurs due to the shape of the constraint surfaces when it moves for a certain finite interval.
Depending on the shape of the constraint surface, it could transit to NC (M=3, D=0, C=0) or PC (M=2, D=0, C=1) state. See Figure 8. The TR-NC transition occurs when the edges of the two constraint surfaces are at the same position, and the moving cube simultaneously loses contact with these faces.
@@
This NC (M=3, D=0, C=0) - TR (M=2, D=0, C=1) is the reverse skill of the earlier TR-NC, and for example, a skill such as placing a book in the air in the space between two books would cause the transition from one maintenance dimension to one constraint dimension. Along this dimension, this skill requires the use of vision, as there is no surface contact before the transition.
@@
For the dimension for which the maintenance d-state changes to the constraint d-state, the normal direction of the book in the above example, by applying B9, visual information must be used. Edge information orthogonal to this dimension must be extracted using a visual sensor.
@@ 3.4.3 TR-PC and PC-TR
In the remaining dimension, the constraint d-state transits to the detachment d-state. In the above example, before the transition, the cube is constrained to the top and the bottom surfaces, and after the transition, it emerges onto the surface following the bottom surface. In this direction, B7 can be applied. Before the transition, the motion is adjusted to maintain the constraint d-state. After the transition the motion is adjusted to keep the detachment d-state. However, since detecting this transition is usually difficult, in B7, the control for the detachment d-state is performed throughout the transition. In other words,
@@
Let us consider a cube sliding on a table into a gap between the table surface and a parallel upper wall. A state transition occurs from PC state (M=2, D=1, C=0) on the table to TR state (M=2, D=0, C=1) in the gap. This sliding motion automatically causes such transition and if this cube collides the top wall due to the difference in size, the skill is fundamentally infeasible due to the difference between the sizes of the object (the cube) and the environment (the walls).
@@ 3.5 Other interstate transitions
The similar discussions can be applied for other interstate transitions. For these transitions, only brief descriptions and results are listed below. Detailed derivations of the reward functions are given in Appendix A.
@@ 3.5.1 PR-NC and NC-PR
The PR (M=1, D=0, C=2) - NC (M=3, D=0, C=0) skill involves such as completely pulling a peg out of the hole, while NC-PR involves the opposite skill of swiftly inserting a peg into a hole (See Figure 11). Since there is no physical contact with the environment before the transition, it is necessary to determine the hole position using visual sensors.
@@ 3.5.2 PR-PC and PC-PR
The PR (M=1, D=0, C=2) - PC (M=2, D=1, C=0) transition occurs, for example, when pulling a peg inside a hole, of which a part of the side is extended externally. Thus, when the peg comes out, contact with the extended surface persists, leading to a detachment d-state instead of a maintenance d-state in that direction unlike the PR-NC skill. See Figure 12.
@@ 3.5.3 PR-TR and TR-PR
PR (M=1, D=0, C=2) - TR (M=2, D=0, C=1) is a skill required in situations where, using the previous example, a pair of opposite surfaces extend outside the hole, and despite the transition from the constraint d-state to the maintenance d-state occurs in the other orthogonal direction, this direction remains in the constraint d-state. See Figure 13.
@@ 3.5.4 PR-OT and OT-PR
PR (M=1, D=0, C=2) - OT (M=1, D=1, C=1) is a skill required in situations where surface contact continues in three directions. See Figure 14.
@@ 3.5.5 OP-PR and PR-OP
The OP (M=0, D=1, C=2) state is like a peg that has reached the bottom of a hole and can move only in the half direction away from the bottom, with two orthogonal directions to this direction constrained, while the PR (M=1, D=0, C=2) state is like a peg that remains in the middle of a hole and can move along the hole in both directions. In the previous paper, we name OP-PR and PR-OP as PTG51 (drawer-opening) and PTG53 (drawer-closing). We can obtain those reward functions as shown in Figure 15.
@@ 3.5.6 OT-NC and NC-OT
OT (M=1, D=1, C=1) - NC (M=3, D=0, C=0) is a skill that involves pulling an object along the detachment surface, not in the normal direction of the detachment surface, when the object is surrounded by a pair of opposing directions and another direction, as shown in Figure 16. In the reverse skill, NC-OT, visual feedback is required.
@@ 3.5.7 OT-PC and PC-OT
The previous OT-NC transition occurs while an object in the center is in motion with maintaining contact with the surrounding environment, the contact ends simultaneously due to the shape of the environment. On the other hand, the OT (M=1, D=1, C=1)-PC (M=2, D=1, C=0) transition occurs, while the object is moving, contact with the environment continues in one direction. In this OT - PC transition, the control differs depending on whether the contact with original detachment surface continues or the contact with a part of the surfaces that created the constraint d-state persists: (a) the detachment d-state continues and (b) the constraint d-state transitions to a detachment d-state. See the details in Appendix.
@@ 3.5.8 OT-TR and TR-OT
Regarding the transition of OT (M=1, D=1, C=1) - TR (M=2, D=0, C=1), there are two cases: (a) motion to break the face contact of the detachment surface and (b) motion to maintain the face contact to the detachment surface. Also see the details in Appendix.
@@ 3.6.1 NC-NC
The skill of the robot corresponding to NC (M=3, D=0, C=0) - NC is a Bring skill and PTG12 in Ikeuchi et al. (2018). It is the unconstrained motion to bring an object from one position to another.
@@
For the two dimensions orthogonal to the direction of motion, B1 is also applicable since the maintenance d-state is maintained.
@@
  Reward NC-NC (PTG12 (Bring) skill)
@@ 3.6.2 PC-PC
Putting these together, we obtain (See Figure 20):
  Reward PC1-PC1 (STG2 (Wipe) skill)
@@
Figure 20: PC1-PC1 skills
@@
    if T = goal-t AND F-s > delta-zero,
@@
    if S = goal-s AND T = goal-t AND
       F+s < delta-zero, then reward
@@
In the PC2 (M=1, D=2, C=0), a one-dimensional maintenance d-state exists. Along this dimension, a motion to maintain the state is possible. Thus, A1 is applicable.
@@
Concerning the two-dimensions orthogonal to the motion, the detachment d-state is maintained with keeping contact to the environment surfaces. B2 is applicable.
@@
PC2 (M = 1, D = 2, C = 0) has a one maintenance dimension that allows motion while maintaining contact with two different surfaces. This dimension corresponds to the base of the crescent-shaped cone of a Gaussian sphere. When the object collides with a third surface during motion along this maintenance dimension, the transition from PC2 to PCN occurs.
@@
In the two dimensions orthogonal to the motion direction, the detachment state is maintained, so B2 is applicable:
@@ 3.6.4 OT-OT
In OT (M=1, D=1, C=1) state, there are two sub-classes: OT1 (M=1, D=1, C=1) and OT2 (M=0, D=2, C=1). Similar arguments to the case of PC can be applied to transitions between these sub-classes. Reward functions for these skills are given in Figure 25. For detailed derivation, please refer to Appendix B.
@@ 3.7 Rotational skills
We will take an example, the OR (M = 0, D = 1, C = 2) - RV (M = 1, D = 0, C = 2) transition, corresponding to a door-opening skill, which appears particularly frequently. The OR state corresponds to a state where the door is closed, and rotation is possible only in the direction of detachment (i.e., opening direction). The transition in the direction of rotation is from the detachment d-state to maintenance d-state, and A3 can be applied. In other words, the drag force in the direction of rotation, F+s, is eliminated, and the skill concludes when a certain predetermined rotation angle, goal-s, given by the demonstration, is reached.
@@
The two degrees of freedom concerning the direction of the rotation axis are both constrained, resulting in B3. In other words, attempting to rotate the rotation axis forcibly will generate drag forces, F-t and F-u. Therefore, it is necessary to maintain these forces below a certain threshold.
@@
  Reward OR-RV (PTG51 (Door-open) skill)
@@
  Reward RV-OR (PTG53 (Door-close) skill)
@@
  Reward RV-RV (PTG52 (Door-adjust) skill)
@@
4 Implementation of skill library
4.1 Current skill library
The manipulation skill library currently consists of a commonly used set of skills, including
  • PC-NC-a (PTG11, Pick) skill
  • NC-NC (PTG12, Bring) skill
  • NC-PC-a (PTG13, Place) skill
  • OP-PR (PTG31, Drawer-open) skill
  • PR-PR (PTG32, Drawer-adjust) skill
  • PR-OP (PTG33, Drawer-close) skill
  • OR-RV (PTG51, Door-open) skill
  • RV-RV (PTG52, Door-adjust) skill
  • RV-OR (PTG53, Door-close) skill
  • PC1-PC1 (STG2, Wipe) skill
@@ 4.1 Current skill library
In order to build an end-to-end system, grasping skills are also necessary. The grasping skills described in Saito et al. (2022) have been implemented for Shadow-hand and Fetch Parallel-gripper in the grasp skill library. Currently, this library includes:
  • Active-force grasp skill
  • Passive-force grasp skill
  • Lazy-closure grasp skill
The execution of a skill, in this case the OP-PR skill, is illustrated in Figure 26. First, a sequence of task models is generated from the human demonstration. Each task model has the skill parameters to execute that task based on the demonstration. In the OP-PR task, the direction of pushing the drawer is stored as Axis direction.
Figure 26: Task and skill
@@ 4.2 Reinforcement learning environment
Many skills require prior learning to have policies to adjust motion directions based on the force feedback with reward functions. For training those skills, we developed a reinforcement learning environment that parallelized the PPO algorithm in Stable Baselines3222https://stable-baselines3.readthedocs.io/en/master/.
To build the learning environment, we used PyBullet [^3] https://pybullet.org/wordpress/ as a simulator and obtained the necessary policies for each skill through reinforcement learning. For example, we set up the environment shown in Figure 27 in the simulator that satisfies the surface contact state. Gravity is set to 0 to keep the environment as simple as possible. We assumed that the object is already grasped by a hand and they are integrated into one unit. Since the estimation of the normal direction of the contact surface may contain some errors, domain randomization was used to add the errors in the training. The contact was assumed to be represented by an elastic body, and hardware compliance was simulated by setting the spring and damper coefficients.
The control side of each skill is described by a force relationship. It is assumed that the actual robot is equipped with a 6-axis force sensor. However, due to the difference in physical parameters between the simulator and the actual machine, it is difficult to simulate the magnitude of the force accurately. Therefore, to ensure that the policy obtained in the simulation can be implemented on the actual robot, we used the unit vector of the force, $\mathbf{f}$ instead of the original value as the states for the reinforcement learning.
$\mathbf{f}_{n}$ $=$ $\mathbf{f}/|\mathbf{f}|$ (2)
The values of the force sensor shall be represented by converting them to the world coordinate system using the position and posture of the object obtained on the simulator and the FK of the actual machine.
@@
$f_{desc}=\lfloor|\mathbf{f}|/f_{step}\rfloor$ (3)
Here, $f_{step}$ is an arbitrary constant that determines the degree of discrimination, and different values are used for simulator and the actual machine to bridge the difference between the simulator and the actual machines.
4.3 Skills
4.3.1 PC-NC-a (pick), NC-NC (bring), NC-PC-a (place) skills
@@
The PC-NC-a skill can be described by the following control rule.
@@
Since the position of the hand before the execution of the skill is known from the end value of the previous skill, the values of goal-s, goal-t, ant goal-u can be calculated using the value in the EDC slot. The skill ends when the target position is reached. The NC-NC skill is implemented in much the same way.
The NC-PC-a skill, also called PTG13 (Place), is also almost the same as the PC-NC-a skill except for the termination condition. The end position obtained from the demonstration includes an observation error, thus, the control algorithm,
@@
Here we show the results of the execution using a real robot to confirm that the implemented skills are working correctly. Figure 29 shows the NC-PC-a skill execution. Figure 30 shows the change in force sensor values. To remove the effect of gravity, the state of the force at the start of the skill is retained and the difference from it is calculated; by using the relationship between the force sensor coordinate system in the robot coordinate system obtained from FK, the force values are converted to those in the world coordinate system, showing the change in force in the z-axis direction (vertical upward). It can be confirmed that the force increases due to the drag force generated when the grasping object comes into contact with the surface. 3 [N] was set as delta-zero, so contact was detected at time 12 and this skill was completed correctly.
Figure 29: Execution of NC-PC-a (PTG13,Place) skill
@@
4.3.2 OP-PR (Drawer-open), PR-PR (Drawer adjust), PR-OP (Drawer-close) skills
These skills can be described by the following control rule without the termination condition.

### 20 Mar 2024 v2

@@ 3.4.2 TR-NC and NC-TR
TR (M=2, D=0, C=1) has two maintenance dimensions and one constraint dimension and a typical example of an object in this state is a cube sandwiched between two parallel walls. In this case, the normal direction of the wall is in the constraint dimension. The cube can move in the maintenance directions (i.e., in the directions parallel to the walls). This maintenance motion itself does not cause a state transition of the object in an infinitesimal interval, but the transition occurs due to the shape of the constraint surfaces when it moves for a certain finite interval.
Depending on the shape of the constraint surface, it could transit to NC (M=3, D=0, C=0) or PC (M=2, D=0, C=1) as shown in Figure 8. The TR-NC transition occurs when the edges of the two constraint surfaces are at the same position, and the moving cube simultaneously loses contact with these faces.
@@
This NC (M=3, D=0, C=0) - TR (M=2, D=0, C=1) is the reverse skill of the earlier TR-NC, and for example, a skill such as placing a book in the air into the space between two books would cause the transition from one maintenance dimension to one constraint dimension. Along this dimension, this skill requires the use of vision, as there is no surface contact before the transition.
@@
For the dimension for which the maintenance d-state changes to the constraint d-state (the normal direction of the book in the above example), B9 can be applied and visual information must be used.
@@ 3.4.3 TR-PC and PC-TR
In the remaining dimension, the constraint d-state transits to the detachment d-state. In the above example, before the transition, the cube is constrained to the top and the bottom surfaces, and after the transition, it detaches from the top surface by sliding the bottom surface. In this direction, B7 can be applied.
@@
Let us consider a cube sliding on a table into a gap between the table surface and a parallel upper wall. A state transition occurs from PC (M=2, D=1, C=0) on the table to TR (M=2, D=0, C=1) in the gap. As described above, this sliding motion automatically causes such transition and if this cube collides the top wall due to the difference in size, the skill is fundamentally infeasible due to the difference between the sizes of the object (the cube) and the environment (the walls).
@@ 3.5 Other interstate transitions
The similar discussions can be applied for other interstate transitions. For these transitions, only brief descriptions and results are listed below. Detailed derivations of the reward functions are given in Appendix A and Figures 11 to 18 show their reward functions.
@@ 3.5.1 PR-NC and NC-PR
PR (M=1, D=0, C=2) - NC (M=3, D=0, C=0) involves such as completely pulling a peg out of the hole, while NC-PR involves the opposite skill of swiftly inserting a peg into a hole. Since there is no physical contact with the environment before the transition, it is necessary to determine the hole position using visual sensors.
@@ 3.5.2 PR-PC and PC-PR
The PR (M=1, D=0, C=2) - PC (M=2, D=1, C=0) transition occurs, for example, when pulling a peg inside a hole, of which a part of the side is extended externally. Thus, when the peg comes out, contact with the extended surface persists, leading to a detachment d-state instead of a maintenance d-state in that direction unlike the PR-NC skill.
@@ 3.5.3 PR-TR and TR-PR
PR (M=1, D=0, C=2) - TR (M=2, D=0, C=1) is a skill required in situations where, using the previous example, a pair of opposite surfaces extend outside the hole, and despite the transition from the constraint d-state to the maintenance d-state occurs in the other orthogonal direction, this direction remains in the constraint d-state.
@@ 3.5.4 PR-OT and OT-PR
PR (M=1, D=0, C=2) - OT (M=1, D=1, C=1) is a skill required in situations where surface contact continues in three directions.
@@ 3.5.5 OP-PR and PR-OP
OP (M=0, D=1, C=2) is like a peg that has reached the bottom of a hole and can move only in the half direction away from the bottom, with two orthogonal directions to this direction constrained, while PR (M=1, D=0, C=2) is like a peg that remains in the middle of a hole and can move along the hole in both directions. In the previous paper, we name OP-PR and PR-OP as PTG31 (Drawer-opening) and PTG33 (Drawer-closing).
@@ 3.5.6 OT-NC and NC-OT
OT (M=1, D=1, C=1) - NC (M=3, D=0, C=0) is a skill that involves pulling an object along the detachment surface, not in the normal direction of the detachment surface, when the object is surrounded by a pair of opposing directions and another direction. In the reverse skill, NC-OT, visual feedback is required.
@@ 3.5.7 OT-PC and PC-OT
The previous OT-NC transition occurs while an object in the center is in motion with maintaining contact with the surrounding environment, the contact ends simultaneously due to the shape of the environment. On the other hand, the OT (M=1, D=1, C=1)-PC (M=2, D=1, C=0) transition occurs, while the object is moving, contact with the environment continues in one direction. In this OT - PC transition, the control differs depending on whether the contact with original detachment surface continues or the contact with a part of the surfaces that created the constraint d-state persists: (a) the detachment d-state continues and (b) the constraint d-state transits to the detachment d-state. See the details in Appendix.
@@ 3.5.8 OT-TR and TR-OT
Regarding the transition of OT (M=1, D=1, C=1) to TR (M=2, D=0, C=1), there are two cases: (a) motion to break the face contact of the detachment surface and (b) motion to maintain the face contact to the detachment surface. Also see the details in Appendix.
@@ 3.6.1 NC-NC
The skill of the robot corresponding to NC (M=3, D=0, C=0) - NC is a Bring task and PTG12 in Ikeuchi et al. (2018). It is the unconstrained motion to bring an object from one position to another.
@@
For the two dimensions orthogonal to the motion direction, B1 is also applicable since the maintenance d-state is maintained.
@@
  Reward NC-NC (PTG12 (Bring) task)
@@ 3.6.2 PC-PC
Putting these together, we obtain:
  Reward PC1-PC1 (STG2 (Wipe) task)
@@
Figure 20 shows the summary of the PC1-PC1 skill.
Figure 20: PC1-PC1 skill
@@
    if F-s > delta-zero AND T = goal-t,
@@
    if F+s < delta-zero AND S = goal-s AND
       T = goal-t, then reward
@@
In PC2 (M=1, D=2, C=0), a one-dimensional maintenance d-state exists. Along this dimension, a motion to maintain the state is possible. Thus, A1 is applicable.
@@
Concerning the two dimensions orthogonal to the motion, the detachment d-state is maintained with keeping contact to the environment surfaces. B2 is applicable.
@@
PC2 (M=1, D=2, C=0) has one maintenance dimension that allows motion while maintaining contact with two different surfaces. This dimension corresponds to the base of the crescent-shaped cone of a Gaussian sphere. When the object collides with a third surface during motion along this maintenance dimension, the transition from PC2 to PCN (M=0, D=3, C=0) occurs.
@@
In the two dimensions orthogonal to the motion direction, the detachment d-state is maintained, so B2 is applicable:
@@ 3.6.4 OT-OT
In the OT state, there are two sub-classes: OT1 (M=1, D=1, C=1) and OT2 (M=0, D=2, C=1). Similar arguments to the case of the PC state can be applied to transitions between these sub-classes. Reward functions for these skills are given in Figure 25. For detailed derivation, please refer to Appendix B.
@@ 3.7 Rotational skills
We will take an example, the OR (M = 0, D = 1, C = 2) - RV (M = 1, D = 0, C = 2) transition, corresponding to a Door-opening task, which appears particularly frequently. OR corresponds to a state where the door is closed, and rotation is possible only in the direction of detachment (i.e., opening direction). The transition in the direction of rotation is from the detachment d-state to the maintenance d-state, and A3 can be applied. In other words, the drag force in the direction of rotation, F+s, is eliminated, and the skill concludes when a certain predetermined rotation angle, goal-s, given by the demonstration, is reached.
@@
The two other dimensions are both constrained, resulting in B3. In other words, attempting to rotate with infeasible displacement forcibly will generate drag forces, F-t and F-u. Therefore, it is necessary to maintain these forces below a certain threshold.
@@
  Reward OR-RV (PTG51 (Door-open) task)
@@
  Reward RV-OR (PTG53 (Door-close) task)
@@
  Reward RV-RV (PTG52 (Door-adjust) task)
@@
4 Implementation of skill-agent library
4.1 Current skill-agent library
The manipulation skill-agent library currently consists of a commonly used set of skill agents, including
  • PC-NC-a (PTG11, Pick) skill agent
  • NC-NC (PTG12, Bring) skill agent
  • NC-PC-a (PTG13, Place) skill agent
  • OP-PR (PTG31, Drawer-open) skill agent
  • PR-PR (PTG32, Drawer-adjust) skill agent
  • PR-OP (PTG33, Drawer-close) skill agent
  • OR-RV (PTG51, Door-open) skill agent
  • RV-RV (PTG52, Door-adjust) skill agent
  • RV-OR (PTG53, Door-close) skill agent
  • PC1-PC1 (STG2, Wipe) skill agent
@@ 4.1 Current skill-agent library
In order to build an end-to-end system, grasping skill agents are also necessary. The grasping skill agents described in Saito et al. (2022) have been implemented for Shadow-hand and Fetch Parallel-gripper in the grasp skill-agent library. Currently, this library includes:
  • Active-force grasp skill agent
  • Passive-force grasp skill agent
  • Lazy-closure grasp skill agent
The execution of a skill agent, in the case of the OP-PR task, is illustrated in Figure 26. First, a sequence of task models is generated from the human demonstration. Each task model has the skill parameters to execute that task based on the demonstration. In the OP-PR task, the direction of pushing the drawer is stored as Axis direction.
Figure 26: Task and skill agent
@@ 4.2 Reinforcement learning environment
Many skill agents require prior learning policies to adjust motion directions based on the force feedback with reward functions. For training those skill agents, we developed a reinforcement learning environment that parallelized the PPO algorithm in Stable Baselines3222https://stable-baselines3.readthedocs.io/en/master/.
To build the learning environment, we used PyBullet [^3] https://pybullet.org/wordpress/ as a simulator and obtained the necessary policies for each skill agent through reinforcement learning. For example, we set up the environment shown in Figure 27 in the simulator that satisfies the surface contact state. Gravity is set to 0 to keep the environment as simple as possible. We assumed that the object is already grasped by a hand and they are integrated into one unit. Since the estimation of the normal direction of the contact surface may contain some errors, domain randomization was used to add the errors in the training. The contact was assumed to be represented by an elastic body, and hardware compliance was simulated by setting the spring and damper coefficients.
The control side of each skill agent is described by a force relationship. It is assumed that the actual robot is equipped with a 6-axis force sensor. However, due to the difference in physical parameters between the simulator and the actual machine, it is difficult to simulate the magnitude of the force accurately. Therefore, to ensure that the policy obtained in the simulation can be implemented on the actual robot, we used the unit vector of the force $\mathbf{f}$ instead of the original value as the states for the reinforcement learning.
$\mathbf{f}_{n}$ $=$ $\mathbf{f}/|\mathbf{f}|.$ (2)
The values of the force sensor shall be represented by converting them to the world coordinate system using the posture of the object obtained on the simulator and the FK of the actual machine.
@@
$f_{desc}=\lfloor|\mathbf{f}|/f_{step}\rfloor.$ (3)
Here, $f_{step}$ is an arbitrary constant that determines the degree of discretization, and different values are used for simulator and the actual machine to bridge the gap between the simulator and the actual machines.
4.3 Skill agents
4.3.1 PC-NC-a (Pick), NC-NC (Bring), NC-PC-a (Place) skill agents
@@
The PC-NC-a skill agent can be described by the following control rule.
@@
Since the position of the hand before the execution of the skill is known from the end position of the previous skill, the values of goal-s, goal-t, ant goal-u can be calculated using the value in the EDC slot. The skill ends when the target position is reached. The NC-NC skill agent is implemented in the same way.
The NC-PC-a skill agent, also called PTG13 (Place), is also almost the same as the PC-NC-a skill agent except for the termination condition. The end position obtained from the demonstration includes an observation error, thus, the control algorithm,
@@
Here we show the results of the execution using a real robot to confirm if the implemented skill agents are working correctly. Figure 29 shows the execution of the NC-PC-a skill agent. Figure 30 shows the change in force sensor values. To remove the effect of gravity, the force at the start of the skill is retained and the difference from it is calculated; by using the relationship between the force sensor coordinate system in the robot coordinate system obtained from FK, the force values are converted to those in the world coordinate system. Figure 30 shows the changes in the z-axis direction (vertical upward). It can be confirmed that the force increases due to the drag force generated when the grasping object comes into contact with the surface. Delta-zero was set to 3 [N]. The contact was detected at time 12 and this skill was completed correctly.
Figure 29: Execution of NC-PC-a (PTG13, Place) skill agent
@@
4.3.2 OP-PR (Drawer-open), PR-PR (Drawer-adjust), PR-OP (Drawer-close) skill agents
These skill agents can be described by the following control rule without the termination condition.